# Scalable Multi-Task Learning for Particle Collision Event Reconstruction with Heterogeneous Graph Neural Networks

### v2

@@ 4.3 Reconstruction performance
Table 2 compares the reconstruction performance of the DFEI with the proposed GNN and HGNN models for inclusive beauty-hadron collision events. All models significantly outperform the DFEI based on the percentage of perfect hierarchy and complete reconstruction with a much lower occurrence of not isolated decays and a similar level of partial reconstruction. For instance, the HGNN architecture, which slightly outperforms the other models, has a perfect reconstruction 4.8 times that of the DFEI.
@@
When early pruning is used to accelerate the inference time, there can be a significant drop in performance if the pruning cuts are too tight. This is illustrated in Figure 4, which shows the trade-off between the reconstruction performance and the average GPU inference time for track multiplicities above 300 using the HGNN and WHGNN models. A gradual tightening of pruning selections in layers 1-3 are applied ranging from $\hat{y}^{e_{\rm tr}}_{1-3}>10^{-7}$ to $\hat{y}^{e_{\rm tr}}_{1-3}>10^{-2}$ in multiples of 10. The WHGNN exhibits greater invariance in performance with pruning, which allows for smaller inference times with minimal reconstruction performance loss.
Figure 4: Relationship between the average GPU timing (ms) for track multiplicity above 300 and percentage of each reconstruction category for HGNN (left) and WHGNN (right) models with no early pruning (the points with the highest average GPU timing) and edge pruning in layers 1-3 for selections ranging from $\hat{y}^{e_{\rm tr}}_{1-3}>10^{-7}$ to $\hat{y}^{e_{\rm tr}}_{1-3}>10^{-2}$ in multiples of 10.
The reconstruction performance can be further quantified for several exclusive beauty-hadron decays with different decay topologies. Table 3 compares the performance of the DFEI model (DF) and HGNN architecture (H1) for a variety of specific decay modes. While the HGNN often outperforms the DFEI in terms of perfect and complete reconstruction categories and has a smaller fraction of not isolated reconstruction, it suffers from larger levels of partial reconstruction. For comparison, an equivalent HGNN (H2) was trained with the addition of 3000 training samples split uniformly between the decay modes, which are shown in the upper half of Table 3. H2 achieves a much higher percentage of perfect hierarchy, while maintaining a 1-2% level of partial reconstruction. Moreover, H2 generalises to similar exclusive decays not seen during training, as demonstrated in the lower half of Table 3.
@@
Table 3: Comparison of the percentage of each reconstruction category for DFEI (DF) [22] and two HGNN models (H1 and H2) with tight pruning, for several exclusive beauty-hadron decays. H1 was trained using only inclusive beauty-hadron decays while the training of H2 includes a sample of 3000 events split between the decays in the top half the table. We consider the following subdecays $D^{+}[K\pi\pi]$, $D^{-}_{s}[KK\pi]$, $J/\psi[\mu\mu]$, $\phi[KK]$, $K^{*0}[K\pi]$ and $\Lambda^{+}_{c}[pK\pi]$.
@@ 4.4 PV association
A key additional capability of the HGNN is PV association, which is indicated in the collision event in Figure 5, by the blue connections between tracks and PVs. This is most valuable for displaced tracks, such as those arising from beauty-hadron decays and other long-lived particles. Therefore, we quantified the PV association for all tracks in the event, tracks from the decays of beauty hadrons and the PV-association of whole beauty hadrons. As discussed in Section 3.3, a given track is associated with a PV by selecting the PV-track edge for that track with the highest last-layer probability score, $\hat{y}^{e_{\rm pv-tr}}_{8}$. Meanwhile, when evaluating the PV association of the HGNN for beauty hadrons we use the ground truth to identify the beauty-hadron tracks and associate the beauty hadron with the PV with which the majority of its tracks are associated.
@@
HGNN $\mathcal{L}^{\rm PV}_{\rm BCE}$ pv-tr 97.83 ± 0.02 96.19 ± 0.14 97.60 ± 0.15
HGNN $\mathcal{L}^{\rm PV}_{\rm BCE}$ pv-tr, tr-tr 99.83 ± 0.01 99.74 ± 0.04 99.85 ± 0.04
HGNN $\mathcal{L}^{\rm LCA}_{\rm CE},\mathcal{L}^{\rm PV}_{\rm BCE},\mathcal{L}^{\rm prune}_{\rm BCE}$ pv-tr, tr-tr 99.88 ± 0.01 99.78 ± 0.04 99.85 ± 0.04
@@
Table 4 reports the average PV association accuracy for the conventional “min IP” method, which assigns each track or beauty hadron to the PV with the smallest impact parameter, alongside several HGNN variants, distinguished by their training objectives and edge definitions. Figure 6 illustrates how each method’s PV-track association performance varies with the number of reconstructed PVs. Across all scenarios, the HGNN models markedly surpass the min IP baseline, achieving up to 99.8% association accuracy for tracks originating from beauty hadrons when track–track edges are incorporated. A similar figure for beauty hadrons is shown in  C.
Figure 6: Dependence of the average PV-track association (%) on PV multiplicity for tracks (left) and beauty tracks (right) using various methods.
@@ 5 Discussion
[…]
@@
Although HGNN architectures only deliver a slight gain in beauty-hadron reconstruction performance compared to their GNN counterparts, they provide the extra capability of precise PV association. The architectural advances proposed in this study have the potential to significantly enhance PV association performance at LHCb. This improvement would, in turn, lead to greater sensitivity across the LHCb physics program by reducing the background contributions and enhancing the resolution of key observables. A more accurate PV association would have a particularly strong impact on the precision of measurements involving decays with missing particles, such as neutrinos, by enabling better determination of the B-hadron flight direction.
@@ 10 Data availability statement
The data and code supporting the findings of this study are openly available at the following URLs: https://zenodo.org/records/15584745 and https://github.com/willsutcliffe/scalable mtl hgnn
@@ Appendix B Multi-task hyperparameter optimisation
Appendix C PV association
Figure 9 compares the dependence of the PV association for beauty hadrons between the minimum impact parameter method and the approaches using HGNNs.
Figure 9: Dependence of the average PV association (%) on PV multiplicity for beauty hadrons when using various HGNN variants and the minimum impact parameter method for PV association.

### v3

@@ 4.3 Reconstruction performance
Figure 4: Relationship between the average GPU timing (ms) for events with track multiplicity greater than 300 and percentage of each reconstruction category for HGNN (top) and WHGNN (bottom) models. The left panels show the fractions of Perfect and Complete reconstructions, while the right panels show the fractions of Partial and Not isolated reconstructions. The results are shown for configurations without early pruning (corresponding to the triangular points with the highest average GPU timing) and with edge pruning applied in layers 1-3 for selection thresholds ranging from $\hat{y}^{e_{\rm tr}}_{1-3}>10^{-7}$ to $\hat{y}^{e_{\rm tr}}_{1-3}>10^{-2}$ in factors of 10.
Table 2 compares the reconstruction performance of the DFEI with the proposed GNN and HGNN models for inclusive beauty-hadron collision events. All models significantly outperform the DFEI in terms of the percentage of perfect and complete reconstructions, exhibiting a much lower occurrence of non-isolated decays and a comparable level of partial reconstruction. In particular, the HGNN architecture, which slightly surpasses the other models, achieves a perfect reconstruction rate 4.8 times higher than that of the DFEI. The quoted uncertainties correspond to standard deviations estimated in the same way as for the LCAG class accuracies: for each reconstruction category, the fraction of events in that category is treated as a binomial ratio, giving $\sigma=\sqrt{p(1-p)/N}$ with $p$ the observed fraction and $N$ the total number of events.
When early pruning is used to accelerate the inference time, there can be a significant drop in performance if the pruning cuts are too tight. This is illustrated in Figure 4, which shows the trade-off between the reconstruction performance and the average GPU inference time for track multiplicities above 300 using the HGNN and WHGNN models. A gradual tightening of pruning selections in layers 1-3 are applied ranging from $\hat{y}^{e_{\rm tr}}_{1-3}>10^{-7}$ to $\hat{y}^{e_{\rm tr}}_{1-3}>10^{-2}$ in factors of 10. The WHGNN exhibits greater invariance in performance with pruning, which allows for smaller inference times with minimal reconstruction performance loss.
@@
Table 3: Comparison of the percentage of each reconstruction category for DFEI (DF) [22] and two HGNN models (H1 and H2) with tight pruning, for several exclusive beauty-hadron decays. Both DFEI and H1 were trained only on inclusive beauty-hadron events, while H2 additionally includes a sample of 3000 events split between the decays listed in the upper half of the table. We consider the following subdecays $D^{+}[K\pi\pi]$, $D^{-}_{s}[KK\pi]$, $J/\psi[\mu\mu]$, $\phi[KK]$, $K^{*0}[K\pi]$ and $\Lambda^{+}_{c}[pK\pi]$.
The reconstruction performance can be further quantified for several exclusive beauty-hadron decays with different decay topologies. Table 3 compares the performance of the DFEI model (DF) and HGNN architecture (H1) for a variety of specific decay modes. While the HGNN often outperforms the DFEI in terms of perfect and complete reconstruction categories and has a smaller fraction of not isolated reconstruction, it suffers from larger levels of partial reconstruction. For comparison, an equivalent HGNN (H2) was trained with the addition of 3000 training samples split uniformly between the decay modes, which are shown in the upper half of Table 3. H2 achieves a much higher percentage of perfect reconstruction, while maintaining a 1-2% level of partial reconstruction. Moreover, H2 generalises to similar exclusive decays not seen during training, as demonstrated in the lower half of Table 3.
@@ 4.4 PV association
A key additional capability of the HGNN is PV association, which is indicated in the collision event in Figure 5, by the blue connections between tracks and PVs. This is most valuable for displaced tracks, such as those arising from beauty-hadron decays and other long-lived particles. Therefore, we quantified the PV association for all tracks in the event, tracks from the decays of beauty hadrons and the PV-association of whole beauty hadrons. As discussed in Section 3.3, a given track is associated with a PV by selecting the PV-track edge for that track with the highest last-layer probability score, $\hat{y}^{e_{\rm pv-tr}}_{8}$. Meanwhile, when evaluating the PV association of the HGNN for beauty hadrons we use the ground truth to identify the beauty-hadron decay tracks and associate the beauty hadron with the PV with which the majority of its tracks are associated.
@@
MLP $\mathcal{L}^{\mathrm{P}V}_{\mathrm{B}CE}$ - 96.37 ± 0.03 90.27 ± 0.20 96.35 ± 0.17
HGNN $\mathcal{L}^{\mathrm{P}V}_{\mathrm{B}CE}$ pv-tr 97.83 ± 0.02 96.19 ± 0.14 97.60 ± 0.15
HGNN $\mathcal{L}^{\mathrm{P}V}_{\mathrm{B}CE}$ pv-tr, tr-tr 99.83 ± 0.01 99.74 ± 0.04 99.85 ± 0.04
HGNN $\mathcal{L}^{\mathrm{P}V}_{\mathrm{B}CE},\mathcal{L}^{\mathrm{p}rune}_{\mathrm{B}CE},\mathcal{L}^{\mathrm{L}CA}_{\mathrm{C}E}$ pv-tr, tr-tr 99.88 ± 0.01 99.78 ± 0.04 99.85 ± 0.04
@@
Table 4 reports the average PV association accuracy per event for the conventional “min IP” method, which assigns each track or beauty hadron to the PV with the smallest impact parameter, alongside an MLP baseline and several HGNN variants, distinguished by their training objectives and edge definitions. The MLP, which serves as a non-GNN baseline, is further described in  C. To compute the average PV association accuracy for track / beauty hadron with PV associations, we first calculate, for each event, the PV association accuracy as the ratio of correct associations (numerator) to the total number of associations (denominator). The reported value corresponds to the mean of these per-event accuracies across all events, and the quoted uncertainties represent the standard error of the mean.
The results show that the HGNN models consistently outperform the conventional min IP approach and MLP baseline across all categories. Incorporating both PV-track and track-track edges yields a substantial improvement, increasing the PV association accuracy for beauty-hadron tracks to around 99.7%. This demonstrates that relational information between tracks provides a powerful discriminant for identifying the correct PV, effectively resolving ambiguities that the impact-parameter method cannot distinguish. Furthermore, the inclusion of auxiliary tasks, such as the LCA and pruning objectives, yields a small but consistent additional gain, indicating that joint optimization across related tasks improves the network’s ability to model the full event topology.
Figure 6 and Figure 7 illustrate how the PV association performance of different methods varies with the number of reconstructed PVs. Figure 6 shows the dependence of the average PV-track association accuracy for all tracks (left) and beauty-hadron decay tracks (right), while Figure 7 presents the corresponding results for beauty hadrons. As expected, the association accuracy decreases with increasing PV multiplicity, reflecting the growing ambiguity in assigning tracks or hadrons to the correct PV in more complex events. The conventional min IP method followed by the MLP baseline exhibit the strongest degradation, particularly for beauty-hadron decay tracks, due to their displaced nature arising from the relatively long beauty-hadron lifetime. In contrast, all HGNN variants maintain significantly higher accuracy across the full range of PV multiplicities, demonstrating improved robustness in high-occupancy environments. Among the HGNN configurations, the models incorporating both track-track and PV-track edges show the best overall performance, with the multi-task variant providing a modest additional gain.
Figure 6: Dependence of the average PV-track association (%) on PV multiplicity for tracks (left) and beauty tracks (right) using various methods for PV association.
Figure 7: Dependence of the average PV association (%) on PV multiplicity for beauty hadrons using various methods for PV association.
@@ 5 Discussion
The GNN and HGNN architectures with integrated pruning demonstrate substantially improved reconstruction performance over DFEI for inclusive beauty-hadron decays. This improvement was driven by the effective background isolation achieved via tight pruning selections and improved LCAG classification. Although the initial training of the HGNN (H1 in Table 3) showed a tendency to prune signal particles in rare exclusive hadronic decays ($\sim$$10^{-5}$ frequency or below [56]), a potential issue for trigger applications, this was effectively addressed. By incorporating a subset of these exclusive decays into the HGNN training process (H2 in Table 3), perfect reconstruction rates exceeding 90% were achieved for these challenging modes, with improvements generalising to similar decays. This highlights the importance of tailored training data to optimise performance on rare topologies. The resulting overall gains in performance would have a major impact on both data acquisition and offline analysis, enabling the highly efficient retention of diverse beauty-hadron decays with minimal background.
@@
In addition to providing substantial efficiency gains over DFEI, the HGNN architecture offers the important advantage of enabling precise PV association. The architectural advances proposed in this study have the potential to significantly enhance PV association performance at LHCb. This improvement would, in turn, lead to greater sensitivity across the LHCb physics program by reducing the background contributions and enhancing the resolution of key observables. A more accurate PV association would have a particularly strong impact on the precision of measurements involving decays with missing particles, such as neutrinos, by enabling better determination of the B-hadron flight direction.
@@ 10 Data availability statement
The data and code supporting the findings of this study are openly available at the following URLs: https://zenodo.org/records/15584745 and https://github.com/willsutcliffe/scalable_mtl_hgnn
@@ Appendix B Multi-task hyperparameter optimisation
Appendix C MLP for PV association
The MLP baseline is a feed-forward neural network with five fully connected layers of 128 hidden units and ReLU activations, trained with a binary cross-entropy (BCE) loss to predict whether a given PV-track or PV-beauty-hadron pair corresponds to a correct association.
For PV-track pairs, the input features are equivalent to the node and edge features used by the HGNN. Specifically, for a given PV-track edge, the inputs include the track and PV node features described in Section 3.4, together with the edge feature, namely the impact parameter.
For PV-beauty-hadron pairs, additional feature engineering is applied. The inputs comprise the summed momentum of the tracks originating from the beauty hadron, the decay origin of the beauty hadron, the PV position coordinates, the impact parameter (IP) between the summed momentum of the tracks and the PV, and two boolean indicators specifying whether the PV corresponds to the minimum-IP or second from minumum-IP PV on an event level.
When applied to an event, the network outputs a probability score for the association of each track or beauty hadron to each PV candidate. The associated PV is taken as the one with the highest probability. Training uses the same dataset splits as for the HGNNs.